OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 3x – 4 A. (11, 5) and (2, 2) B. (5, 11) and (3, 2) C. (5, 11) and (2, 3) D. (5, 11) and (2, 2)

jimthompson5910 (jim_thompson5910):

Let's check (11,5) (x,y) = (11,5) which means x = 11 and y = 5 y = 3x – 4 5 = 3*11 – 4 ... replace x with 11, replace y with 5 5 = 33 - 4 5 = 29 The equation 5 = 29 is false. Therefore, (11,5) is NOT a solution to the equation y = 3x-4

jimthompson5910 (jim_thompson5910):

Since (11,5) is NOT a solution, we can rule out choice A.

jimthompson5910 (jim_thompson5910):

You need to check both points, for each choice, to see which make the equation true. Both need to make the equation true.

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome